(D² - 2DD' + D'²)z = x²y²ex+y


1
Expert's answer
2021-06-08T04:39:22-0400

"(D-D')^2z=0"


General solution:

C.F+P.I


C.F.:

"\phi_1(y+x)+x\phi_2(y+x)"


P.I.:

"\frac{1}{(D-D')(D-D')}(x²y²e^x+y)"


"\frac{1}{D-D'}(x²y²e^x+y)=\int(x²y²e^x+y)dx=y^2(x^2-2x+2)e^x+xy"


"\frac{1}{D-D'}(y^2(x^2-2x+2)e^x+xy)=\int(y^2(x^2-2x+2)e^x+xy)dx="


"=y^2(x^2-4x+6)e^x+x^2y\/2"